OpenStudy (anonymous):

integral of (-cosx)/(cosx+sinx) from 0 to pi/2 .. !!! Please explain me how to do this the easiest way =)

OpenStudy (anonymous):

There's an identity in there somewhere..

OpenStudy (wasiqss):

try do rationalization.

OpenStudy (anonymous):

what do you mean ? i don't think we learned it yet ..

OpenStudy (anonymous):

Disregard.

OpenStudy (wasiqss):

lol you did it wrong

OpenStudy (anonymous):

just for checking u doin integral for [0 to pi/2] or [-pi/2 to pi/2]?

OpenStudy (anonymous):

i will tell you, the answer should be pi/4 and i know that somehow you can write the integral as 2*integral of (cosx +sinx)/(cosx+sinx) but i don't understand why ... =(

OpenStudy (anonymous):

from 0 !!! sorry !

OpenStudy (wasiqss):

cause the differential of denominator is present in in numerator

OpenStudy (anonymous):

ok so what we have is\[\int_{0}^{\pi/2} \frac{\cos x}{\cos x+\sin x} \text{d}x\]

OpenStudy (wasiqss):

yup mukushla

OpenStudy (wasiqss):

i got it!

OpenStudy (wasiqss):

take cos x common from both numerator and denominator

OpenStudy (wasiqss):

i guess so it will be easier that way cause the final eq will be \[\int\limits_{?}^{?}1/(1+tanx)\]

OpenStudy (anonymous):

\[I=\int_{0}^{\pi/2} \frac{\cos x}{\cos x+\sin x} \text{d}x\]TT will show us that with sub \(x=\frac{\pi}{2}-x\) we have \[I=\int_{0}^{\pi/2} \frac{\sin x}{\cos x+\sin x} \text{d}x\]

OpenStudy (turingtest):

haha no I won't :)

OpenStudy (anonymous):

lol

OpenStudy (turingtest):

so you did \[u=\pi/2-x\]?

OpenStudy (anonymous):

hmmm wait a minute .. you should replace the limits as well no ?

OpenStudy (turingtest):

that's what I was gonna say

OpenStudy (turingtest):

\[-\int_{\pi/2}^0{\sin u\over\sin u+\cos u}du\]?

OpenStudy (turingtest):

which is what you wrote, ic :P

OpenStudy (anonymous):

and\[-\int_{\pi/2}^0{\sin u\over\sin u+\cos u}du=\int_{0}^{\pi/2}{\sin u\over\sin u+\cos u}du\]

OpenStudy (turingtest):

right, I get it now...

OpenStudy (turingtest):

that integral seems no easier to me though

OpenStudy (anonymous):

how can i continue? i can't add them together because of the different integrands =\

OpenStudy (anonymous):

\[I=\int_{0}^{\pi/2}{\cos x\over\sin x+\cos x}dx=\int_{0}^{\pi/2}{\sin x\over\sin x+\cos x}dx\]\[2I=\int_{0}^{\pi/2}{\sin x+\cos x\over\sin x+\cos x}dx=\int_{0}^{\pi/2}1dx\]

OpenStudy (anonymous):

ohhh i get it !! thanks a-lot to both of you !!!! =)

OpenStudy (turingtest):

that is sweet dude!

OpenStudy (anonymous):

:)